OpenStudy (anonymous):

exact value of sin 5 times pie over 6

OpenStudy (anonymous):

sin (a) = sin (pi - a) Find the value of "a" where "pi - a" = 5pi/6

OpenStudy (anonymous):

whats the exact answerr

OpenStudy (anonymous):

Are you able to find that value of "a" or do you need further help?

OpenStudy (anonymous):

i just need the ecaxt answer

OpenStudy (anonymous):

From my first post, where I have: pi - a = 5pi/6 to solve for "a", you add "a" to both sides and subtract 5pi/6 from both sides. This will give you an identity value for the sine that should be memorized. It will be one of the "special" values when deriving sines.

OpenStudy (anonymous):

Once you get that value of "a", the following will be a very helpful website: http://www.analyzemath.com/Trigonometry_problems/special_triangles.html Also, to convert from radians back to degrees: 180 degrees = "pi" radians You're going to end up with a fraction of "pi" for "a", so use the above to convert to degrees. Then you can use the website for the "special" angles. This is all you need to solve this problem.

OpenStudy (anonymous):

Are you stuck anywhere, @holsmeets ?

OpenStudy (anonymous):

pi - a = 5pi/6 -> a = pi/6 radians = 30 degrees 30 degrees is one of the special angles and eventually you'll have to commit to memory that: sin 30 = 1/2

OpenStudy (anonymous):

All good now, @holsmeets ?

OpenStudy (anonymous):

yes